what is (4x-1) to the 1/2 power = 3 find x
Found 2 solutions by richwmiller, mananth:
Answer by richwmiller(17219)

4x-1=9
4x=8
x=2

Answer by mananth(16949)

what is (4x-1) to the 1/2 power = 3 find x
sqrt(4x-1) =3 take the square of both sides of the eequation
4x-1 = 9
4x=10
x=10/4 = 5/2